OpenStudy (anonymous):

P and Q each stand for whole numbers. P + Q = 1000 P is 150 greater than Q. Calculate the numbers P and Q ?

OpenStudy (anonymous):

Ok so you have P + Q= 1000 and P - Q = 150 Can you solve for P and Q now? Hint: Use either substitution method or elimination method

OpenStudy (anonymous):

Those are simultaneous equations.

Parth (parthkohli):

p = 150 + q So this can be substituted. \( \color{Black}{\Rightarrow q + (150 + q) = 1000 }\)

OpenStudy (lgbasallote):

good explanation @beeqay ^_^

OpenStudy (anonymous):

Thank you, couldn't tag your name. But I loved simultaneous equations when I was young:)

OpenStudy (anonymous):

younger*

OpenStudy (anonymous):

q(1+(150+1)=1000